OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 4 + 2(x − 1) = 5x − (x + 10) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (anonymous):

4+2x-2=5x-x-10 2+2x=4x-10 2+10=4x-2x 12=2x x=6